9.14 A baseball player has a batting average of 0.328. Let X be the number of hits the player gets during 20 times at bat. UsProblem 9.14. The problem assumes that the batters probability of getting a hit stays constant at p=0.328 while he comes upPlease use the central limit theorem.
Answer #1

ANSWER Given that * A baseball player has a batting average p= 0.328. * Let it be the number of hits the player gets during1-6.5 PIXS 1) = P(24 - ) = Plzs 15:52 = P( Z22.65) =0.996 P(x-3) = P(Z < 3-4.0) =P [241.70) = 0.9554 Plxs 6) = P(Z < 6-6.56)